Fermi National Laboratory
 

Accelerator Update for October 11, 2006 - October 13, 2006

 

Wednesday October 11

The day shift began with the Tevatron (TeV) off due to a magnet replacement, with the Recycler, Antiproton Source (Pbar), Linac, Booster, the Main Injector (MI), NuMI, and MiniBooNE all off due to a Preaccelerator H- Source problem. The Collider Detector at Fermilab (CDF) and D-Zero (D0) are both off and in access.

Interlock technicians completed the Booster safety system tests at 9:28 AM.

Operations reported that Linac was turned on at 2:08 PM. A Linac RF station (LRF2) started having problems at 2:12 PM.

Operators had beam to the Booster dump by 6:34 PM.

At 6:59 PM, LRF3 tripped off for the first of two times in eight minutes. It reset slowly.

Operations established beam to MiniBooNE at 8:40 PM, but reported that NuMI beam was falling out of MI. Operations called in an expert.

At 9:33 PM, Operations reported that the I- Source voltage was drooping. Operators increased the pressure.

At 9:41 PM, the I- power supply tripped off. It reset okay.

Operations rebooted the MI low-level RF system at 10:51 PM.

Operations sent protons to the Recycler at 10:51 PM.

Operations rebooted the NuMI BPM nodes and established beam to NuMI at 11 PM.

At 11:02 PM, a Booster power supply (WAPS) tripped off. Operators had the supply back on by 11:15 PM. Operation contacted experts.

Thursday October 12

The midnight shift began with the TeV in access, and with Booster RF vacuum problems holding off beam to MiniBooNE.

At 12:47 AM, Operations reported that a Booster RF station (BRF12) tripped off on a LCW leak indication. Operators bypassed the station, tuned around it, and established beam to MiniBooNE. With the help of a Booster expert, operators found and replaced a bad five-volt fuse.

At 6:52 AM, Operations reported that the I- Source had been losing intensity through out the shift. Operations contacted an expert.

Interlock technicians completed a NuMI safety system test at 11:08 AM.

Interlock technicians completed a MiniBooNE safety system test at 11:22 AM.

A Preaccelerator expert began tuning the I- Source at 3:54 PM.

Operations gave a summary report of the day shift at 4 PM: TeV sector D3 had warmed and the warm up of D2 had started. The Pbar source was off for an Accumulator magnet replacement. BRF6 had failed and was bypassed.

NuMI experts reported that they were able to repair the HV100 LCW leak. They resumed taking beam at 5:23 PM.

Interlock technicians completed a Pbar safety system test around 7 PM.

At 8:22 PM, Operations reported that the I- Source output had dropped a bit, but then leveled out.

A Preaccelerator expert began tuning the I- Source at 9:16 PM.

Operations gave a summary report of the evening shift at 11:59 PM: RF experts completed repairs on BRF6. Preaccelerator experts accessed the I- Source to inspect its cesium boiler. Pbar enclosures were searched and secured for safety system testing.

Friday October 13

The midnight shift began with the TeV in access and with NuMI and MiniBooNE taking beam.

Operators completed a search and secure of the TeV A-E sectors and the Transfer Hall at 5:28 AM for safety system tests.

At 6 AM, Operations reported that two NuMI power supplies (HV101 and HV101A) tripped off and wouldn’t reset. Investigations showed a ground fault indication. Operations called in experts.

Power supply experts reported at 7:33 AM that HV101A had a failed capacitor. The repair would take three to four hours.

Operators prepared the TeV F sector for an access at 7:38 AM for safety system tests.

The Plans for Friday and the Weekend

  1. Replace the Tevatron D3 magnet
  2. Check the Tevatron separator vacuum at sectors D0 and D17 – the TeV separators may require a bake
  3. Replace the Pbar magnet (A6Q14)
  4. Allow the Pelletron experts to conduct necessary maintenance
  5. Complete safety system tests

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/13/2006    email Fermilab